a fluid has mass of 8kg and occupies a volume of 1.5m^3 the pressure of 101.3 kpa. if the specific internal energy is 1800 kj per kg, determine the total enthalpy in mj


1
Expert's answer
2021-11-11T18:05:28-0500

Enthalpy (H)=U+PV

H=U+PV

U=1800kJ/kg

P=101.3Kpa

V=1.5m3

m=8kg

Now internal energy

"U=1800\times8=14400KJ"

Now put value


"H=14400+101.3\times1.5=14.55195MJ"